### v3

@@ Abstract
The recent performance improvements in mixed-integer programming (MIP) went along with a significantly increased complexity of the codes of MIP solvers, which poses challenges in fixing implementation errors. In this paper, we introduce MIP-DD, a solver-independent tool, which is, to the best of our knowledge, the first open-source delta debugger for MIP. Delta debugging is a hypothesis-trial-result approach to isolate the cause of a solver failure. MIP-DD simplifies MIP instances while maintaining the undesired behavior and already supported and motivated fixes for many bugs in the SCIP releases 8.0.4, 8.1.0, and 9.0.0. This translates to an increase of approximately 71.4% more bugfixes than in the same time period before and including some fixes of long-known issues. As we highlight in selected case studies, instances triggering fundamental bugs in SCIP can typically be reduced to a few variables and constraints in less than an hour. This makes it significantly easier to manually trace and check the solution process on the resulting simplified instances. A promising future application of MIP-DD is the analysis of performance bottlenecks, which could very well benefit from simple adversarial instances.
@@
Over the last few decades, solvers for mixed-integer programming (MIP) have witnessed remarkable performance improvements [1], allowing them to solve even large-scale problems despite the NP-hardness of the problem class. However, these performance gains come with increasingly complex algorithms, often posing challenges in identifying and fixing implementation errors. These errors may cause the solver mistakenly claiming unboundedness or infeasibility of the problem, or even returning “optimal solutions” that are indeed infeasible or suboptimal [2, 3, 4, 5].
@@
In this paper, we describe a new open-source tool MIP-DD [6] that can be used to facilitate and accelerate the process of locating implementation errors, commonly called “bugs”, in any complete or heuristic MIP solver. We discuss several successful case studies from the release process for Version 9.0 of the open-source MIP solver SCIP [7].
Let us first review common difficulties arising when debugging MIP solvers. Typical methods to locate bugs in the code are assertions or a debug solution mechanism. Assertions inform the user if certain essential conditions are violated at runtime. A debug solution mechanism informs the user about the invalidation of a feasible reference solution during the solving process. Nevertheless, these methods have their limitations. First, due to computational overhead, assertions and debug solution checks are usually not applied as frequently as needed to directly locate the occurrence of an error. In addition, bugs resulting from limited numerical precision might even pass assertions that incorporate certain tolerances. Second, due to the complex interaction of different solving techniques in MIP solvers, the root cause of an error may occur at an earlier stage than when an assertion fail is detected. Third, when the input instance admits multiple optimal solutions, any techniques that may cut off optimal solutions, such as strong dual reductions [8] or symmetry handling [9], might have to be turned off to guarantee that a particular debug solution is not validly excluded during the solving process, which can unintentionally work around the issue of interest.
@@
While experienced solver developers are trained to deal with these particular difficulties, tracing back the root cause of an error can still become costly or virtually impossible due to the sheer size and runtime of input problems encountered in many bug reports. The goal of MIP-DD is to attack this size and time aspect of debugging by iteratively shrinking the problem size and the covered code base while preserving the reproducibility of the error. The reductions performed by MIP-DD then usually make it easier for a developer to isolate the actual bug on the reduced input and provide a targeted fix also for the original input.
@@
In the literature, this approach is known as delta debugging [10], which is an automated method to isolate the cause of a software failure, driven by a hypothesis-trial-result loop. Delta debugging has been applied successfully in the SAT solver community [11, 12, 13, 4]. To the best of our knowledge, the present work describes the first application of delta debugging to mixed-integer programming and mathematical optimization in general.
@@ 2 Structure of the Delta Debugger
A simplified workflow of MIP-DD can be seen in Figure 1. MIP-DD consists of multiple Modifiers each equipped with a unique strategy to isolate a bug. Each modifier creates a local copy of the problem and modifies it. In the general context of Delta Debugging, we then wish to test the hypothesis that the solver exhibits the same bug also on the modified problem. In order to speed up this process, modifications can be grouped in batches and tested together as one hypothesis. If the hypothesis is confirmed, i.e., the bug is reproduced, then all modifications are applied to the globally stored problem; otherwise, the modification is reverted.
In the following sections, we explain further details of the implementation. In Section 2.1, we first present the modifiers and their strategies. In Section 2.2, we discuss the API by which the modifiers interact with the solver via a solver interface. This API is able to solve the (modified) problem and needs to evaluate if a bug is reproduced. By this design choice, a broader range of bugs can be detected, see also the example discussed in Section 3.5. In Section 2.3, we motivate the ordering mechanism in which MIP-DD calls modifiers to ensure that no potential modification was missed, in particular if batches are used to speed up the process. Finally, in Section 2.4, we provide some general recommendations for using MIP-DD.
2.1 Modification Modules
MIP-DD consists of nine different modifiers each with a unique strategy to modify the problem or the settings in order to simplify the reproducing instance, reduce the code coverage, or accelerate the solving process to make the subsequent debugging easier. In order to prioritize substantial over superficial reductions, the modifiers are ordered in the following sequence:
@@ 2.1 Modification Modules
Module Constraint deletes a constraint from the problem.
@@
Module Variable fixes a variable to the value of the reference solution.
@@
Module Coefficient deletes a coefficient of a fixed variable and balances the left- and right-hand-side of the constraint .
@@
Module Fixing removes a fixed variable from the problem.
@@
Module Setting switches a parameter of the solver to a user-defined value. These values are defined in the target setting file.
@@
Module Side fixes inequalities to with being the reference solution.
@@
Module Objective sets the objective coefficient of a variable to zero.
@@
Module VarRound rounds objective coefficient and the bounds of a variable to integer values.
@@
Module ConsRound rounds the coefficients of a constraint to integer values. The right-hand and left-hand sides are adjusted accordingly.
An important requirement is that each modifier reduction must preserve the feasibility of the reference solution if it is provided. Therefore, in modifiers VarRound and ConsRound the resulting bounds and sides are relaxed to include the reference solution if necessary. Deleted variables are assumed to be fixed to the respective value of the reference solution. Optimality of the reference solution must not be preserved. Note that for infeasible problems, this approach resembles a heuristic method to identify an irreducible infeasible subset [14, 15, 16].
@@
set to 100, the modifier aggregates 1000 constraints to one batch and tests the removal of all 1000 constraints at once, hence reducing the number of solve invocations to 100.
@@ 2.2 Solver Interface
The communication with the solver occurs via the solver interface, an API that allows easy integration of other MIP solvers. The implementation of the solver interface must report back to the modifiers whether a bug is reproduced. To do so, the solver interface should apply several checks to detect the bugs of interest. In the interface to SCIP, the following checks are applied. It is checked whether
the dual bound of the current problem cuts off the objective value of the reference solution. Under the assumption that all modifiers preserve the feasibility of the reference solution, this is a contradiction to the feasibility of the reference solution. In this case, a suboptimality bug is detected.
a solution returned by the solver is cut off by the problem restriction. In this case, an infeasibility bug is detected. For an unbounded problem, the solver sometimes additionally provides a primal ray, which is then also verified to describe an unbounded improving direction.
the primal bound is better than the evaluated objective value of the optimal solution, i.e., if the best solution is cut off.
whether an unexpected error occurs during the solving process, this is also considered a failure.
All these checks are provided by MIP-DD and can be used for any other MIP solver and can be deactivated by parameter.
@@ 2.3 Workflow
When aggregating modifications in batches, valid modifications may be missed, in particular if only a small subset of the aggregated modifications is responsible for the wrong behavior of the solver. Hence, after the first invocation of a modifier, there might be valid modifications remaining in the current problem making additional calls necessary.
MIP-DD uses a stage-round mechanism to call modifiers multiple times in a different order. This mechanism is similar to the workflow of presolving, see, e.g., [17, 18]. In the -th stage only the modifiers with priority at most are called in sequence. This is repeated over several rounds until no more changes are applied. Then, the stage number is incremented or the algorithm terminates if there are no more modifiers to add to the stage cycle. Hence, if in one round the problem remains the same, the next stage is entered to include an additional modifier. Otherwise, the next round is entered to repeat the process with the same set of modifiers.
@@ 2.4 General Recommendations for Applying MIP-DD
Determine batch size
@@
as the limit of the number of solve invocations per modifier.
In this way, every successful solve invocation will simplify a substantial chunk of the problem at once.
As a rule of thumb, the longer the initial solve takes to reproduce the bug, the smaller should be the setting for parameter nbatches
, in order to decouple the expected total runtime from the original problem size.
The final instance with a possibly reduced runtime could then be further improved with a larger number of batches.
Define limits for solves
@@
Suppress unwanted fails
All kinds of fails are taken seriously by default.
However, it can happen that for example, a dual fail turns into a primal fail during the reduction process.
Although discovering further issues is a beneficial side effect, it is sometimes desired to avoid particular transitions of fail types in order to suppress unresolved known issues or prioritize certain bug fixes.
@@
Negative values are reserved for solver-internal errors and positive values encode issues detected by MIP-DD, representing, e.g., a dual fail by 1, a primal fail by 2, and an objective fails by 3.
Exact solution
Especially for numerically sensitive issues it is important to ensure that the reference solution has feasibility violations as small as possible. To illustrate potential confusion due to an inaccurate reference solution, we consider the simple example
@@ 3 Case Study
A preliminary version of MIP-DD has already been used to track down the majority of the bugs fixed since version 8.0.3 of the open-source solver SCIP [7]. In this section, we present selected examples to showcase different types of errors that can occur during the development of a MIP solver and demonstrate how MIP-DD supports the debugging process.
Virtually all MIP solvers employ floating-point arithmetic. In floating-point arithmetic, a limited amount of memory is allocated to represent a number. Hence, not all numbers can be represented accurately. If more digits are needed than allocated the last digits are omitted leading to a loss in precision. Among further problems, calculation in floating-point arithmetic is neither associative nor distributive meaning that is not necessarily equal to and may yield a different result than . For further information about floating-point arithmetic, we refer to [22].
@@ 3.1 Modifying the Settings to Detect Wrong Calculation of Activities in DualInfer
At first we want to highlight the ability of MIP-DD to identify faulty components by manipulating the settings.
For this we take a look at an instance reported by a SCIP user that has been in the backlog for over three years (SCIP git hash d4c0ad7644
Using the so-called aggressive presolve settings leads to an incorrect solution while running SCIP with default settings resulted in the correct solution.
@@
In fact, the reason for the incorrect solution is that Dual-Infer incorrectly calculated the min-max-residuum for the subset . If applied correctly, the maximal activity is calculated by
@@
the maximum residuum is .
@@
The reason is an index shift where instead of and always and were applied.
@@ 3.2 Fixing Invalid Separation of Interior Solutions
In this section we want to analyze a bug appeared on the MIPLIB [23] instance rococoC10-001000
@@
For this problem with minimal value 11460 a solution with value 23706 was declared optimal after 5776 LP iterations.
Applying delta debugging with disabled separators, heuristics, and presolvers as target settings, primal fails suppressed by using as passcodes
, and 1,000 batches, led to the small integer program
in less than an hour. It has the same optimal value as the original problem while a solution of value 12087 was claimed optimal already after 4 LP iterations. The only remaining non-disabling settings are presolving/maxrounds = -1, constraints/linear/maxprerounds = -1, separating/closecuts/freq = 0, and separating/intobj/freq = 0. This means that apart from some presolving only separators closecuts and intobj are required to reproduce this issue. As closecuts is known as a meta separator, these settings indicated that there might be an issue in separator intobj when called by separator closecuts [24, 25, 26]. Separator closecuts determined an interior point of the LP relaxation which is handed to intobj to be separated. Separator intobj relies on the integrality of the objective function, which is exploited by cutting off a solution with a fractional objective value. This is done by rounding this value up and using it as left-hand side of the objective row. The approach implicitly assumes that there is no feasible solution with a smaller objective value than the infeasible solution to be separated since all undiscovered better solutions would otherwise be cut off. However, for the arbitrary LP-interior points computed by separator closecuts, this does not hold in general. It turned out that the value of the separated solution is approximately 11646.03 indeed exceeding the optimal value 11460. Restricting intobj to optimal solutions of LP relaxations finally resolved this issue.
During preliminary experiments, we observed on this issue that using default settings as target settings, results in a larger final instance with 1528 variables and 301 constraints. The runtime was even comparable to the previously presented setup although disabling plugins usually leads to larger solving times of the single solves. An explanation for this behaviour might be given by the complexity of the solver under default settings which under certain circumstances requires a more complex external problem to reproduce the failing internal state. Nevertheless, targeting default settings is desirable for the generation of lightweight but flexible regular tests. And for other issues, also in this case, comparably small reproducing instances could be found.

### v5

@@ Abstract
The recent performance improvements in mixed-integer programming (MIP) have been accompanied by a significantly increased complexity of the codes of MIP solvers, which poses challenges in fixing implementation errors. In this paper, we introduce MIP-DD, a solver-independent tool, which to the best of our knowledge is the first open-source delta debugger for MIP. Delta debugging is a hypothesis-trial-result approach to isolate the cause of a solver failure. MIP-DD simplifies MIP instances while maintaining the undesired behavior. Preliminary versions already supported and motivated fixes for many bugs in the SCIP releases 8.0.1 to 8.1.1. In these versions, MIP-DD successfully contributed to 24 out of all 51 documented MIP-related bugfixes even for some long-known issues. In selected case studies we highlight that instances triggering fundamental bugs in SCIP can typically be reduced to a few variables and constraints in less than an hour. This makes it significantly easier to manually trace and check the solution process on the resulting simplified instances. A promising future application of MIP-DD is the analysis of performance bottlenecks, which could very well benefit from simple adversarial instances.
@@
Over the last few decades, solvers for mixed-integer programming (MIP) have witnessed remarkable performance improvements [1], allowing them to solve even large-scale problems despite the NP-hardness of the problem class. However, these performance gains come with increasingly complex algorithms. For instance, in the last 10 years, the code base of SCIP increased by approximately 44.6%, precisely from 703 965 lines of code in SCIP 3.1 to 1 017 718 lines of code in SCIP 9.1 [2]. This rise in complexity increasingly poses challenges in identifying and fixing implementation errors, so-called bugs. Bugs may cause any kind of undesired solver behavior, especially claiming unboundedness or infeasibility falsely, or even returning “optimal solutions” that are actually infeasible or suboptimal [3, 4, 5, 6].
@@
In this paper, we describe a new open-source tool MIP-DD [7] that can be used to facilitate and accelerate the process of locating implementation errors, “bugs”, in any complete or heuristic MIP solver. We discuss several successful case studies from the release process for Version 9.0 of the open-source MIP solver SCIP [2].
Let us first review common techniques used when debugging MIP solvers.111See also https://scipopt.org/doc/html/DEBUG.php. Assertions are boolean expressions that must evaluate to true and are used by programmers to test assumptions during runtime. If an assertion fails, the run is immediately interrupted so that the invalid state can be inspected when the solver is compiled in debug mode and a software debugger is attached. A debug solution mechanism allows the developer to enable further assertions which require a provided feasible solution to not be invalidated during the whole solving process. Examples of such invalidations include pruning the node containing the debug solution or the application of a cut that renders the solution infeasible.
Despite their usefulness, these methods have their limitations. First, due to computational overhead, assertions and debug solution checks are not present as frequently as necessary in order to pinpoint the exact moment an error occurs. Second, bugs resulting from limited numerical precision might even pass assertions that incorporate certain tolerances. Third, techniques in the MIP solvers, such as strong dual reductions [8] or symmetry handling [9], are allowed to cut off optimal solutions as long as at least one of them remains feasible. For this reason, a failing debug solution check on some optimal solution does not automatically imply a solver error, which complicates efforts to investigate the issue of interest.
@@
While experienced solver developers are trained to deal with these particular difficulties, tracing back the root cause of an error can still become costly or virtually impossible due to the sheer size and runtime of input problems encountered in many bug reports. The goal of MIP-DD is to attack this size and time aspect of debugging by iteratively shrinking the problem size and the covered code base while preserving the reproducibility of the error. The reductions performed by MIP-DD make it easier for a developer to isolate the actual bug on the reduced input and provide a targeted fix also for the original input.
@@
In the literature, this approach is known as delta debugging [10], which is an automated method to isolate the cause of a software failure, driven by a hypothesis-trial-result loop. Delta debugging has been applied successfully in the SAT and SMT solver community [11, 12, 13, 14, 5]. To the best of our knowledge, this work describes the first application of delta debugging to mixed-integer programming and mathematical optimization in general.
@@ 2 Structure of the Delta Debugger
A simplified workflow of MIP-DD can be seen in Figure 1. MIP-DD consists of multiple Modifiers each equipped with a unique strategy to isolate a bug. Each modifier creates a local copy of the problem and modifies it. In the general context of Delta Debugging, we then wish to test the hypothesis that the solver exhibits the same bug also on the modified problem. In order to accelerate this process, modifications can be grouped in batches and tested together as one hypothesis. If the hypothesis is confirmed, i.e., the bug is reproduced, then all modifications are applied to the globally stored problem; otherwise, the modification is reverted.
In the following subsections, we explain the implementation in further detail. In Section 2.1, we first present the modifiers and their strategies. In Section 2.2, we discuss the API by which the modifiers interact with the solver via a solver interface. Through this API we invoke the solving of the (modified) problem and subsequent evaluation if a bug is reproduced such that a broad range of bugs can be detected, see also the example discussed in Section 3.5. In Section 2.3, we motivate the ordering mechanism in which MIP-DD calls modifiers to ensure that no potential modification was missed, in particular, if batches are used to speed up the process. Finally, in Section 2.4, we provide some general recommendations for using MIP-DD.
2.1 Modifiers
MIP-DD consists of nine different modifiers each with a unique strategy to modify the problem or the settings in order to simplify the reproducing instance, reduce the code coverage, or accelerate the solving process to make the subsequent debugging easier. Modifiers are selected sequentially in a fixed order. To prioritize substantial reductions over more superficial reductions, the nine modifiers available in MIP-DD 1.0 are ordered as follows:
@@ 2.1 Modifiers
Modifier Constraint deletes a constraint from the problem.
@@
Modifier Variable fixes a variable to the value of the reference solution.
@@
Modifier Coefficient deletes a coefficient of a fixed variable and balances the left- and right-hand-side of the constraint .
@@
Modifier Fixing removes a fixed variable from the problem.
@@
Modifier Setting switches a parameter of the solver to a user-defined value. These values are defined in the target setting file.
@@
Modifier Side fixes inequalities to with being the reference solution.
@@
Modifier Objective sets the objective coefficient of a variable to zero.
@@
Modifier VarRound rounds objective coefficient and the bounds of a variable to integer values to polish the reduced instance. Polishing facilitates debugging and rules out certain numerical difficulties.
@@
Modifier ConsRound rounds the coefficients of a constraint to integer values to polish the reduced instance. The right-hand and left-hand sides are adjusted accordingly.
Each modifier selects its modifications by iterating over the constraints and variables sequentially in the original order stored in the solver.
An important requirement is that each modification must preserve the feasibility of a provided reference solution. The optimality of a reference solution is not required. Therefore, in modifiers VarRound and ConsRound, the resulting bounds and sides are relaxed to include the reference solution if necessary. Deleted variables are assumed to be fixed to the respective value of the reference solution.
Note that the algorithm behind MIP-DD is similar to the algorithm in [15, Algorithm 1] to compute an irreducible infeasible subset (IIS) [16, 15, 17, 18]. In the upcoming MIP-DD 2.1, an option to calculate an IIS based on the solver results is introduced. When using this option, modifiers are restricted to relaxing problem reductions and only detecting infeasibility is considered as “bug”.
@@
set to 100, the modifier aggregates 1000 consecutive constraints to one batch and tests the removal of all these constraints at once, hence reducing the number of solve invocations to 100.
Because MIP solvers exhibit performance variability, most solvers provide the option to specify a random seed in order to trigger different solving paths. Using this feature, one possible extension of MIP-DD could be to test each batch on multiple random seeds and accept the modification either if all of the seeds trigger the bug, or if at least one of the seeds triggers the bug. The former variant could help to only accept modifications under which the bug is reproduced “robustly”; the latter variant could in turn find more acceptable modifications, thereby yielding stronger reductions of the problem.
@@ 2.2 Solver Interface
The communication with the solver happens in the solver interface, a general API that allows simple integration of any MIP solver. The task of the solver interface is to report to the calling modifier whether a bug is reproduced by the current instance. To do so, the solver interface applies several checks to detect the bugs of interest. In the interface to SCIP, the following checks are applied, which determine whether
the dual bound of the current problem cuts off the objective value of the reference solution. This is a contradiction to the feasibility of the reference solution. In this case, a suboptimality bug is detected.
a solution returned by the solver is cut off by the problem restriction. In this case, an infeasibility bug is detected. For an unbounded problem, the solver may additionally provide a primal ray, which is then also verified to describe an unbounded improving direction.
the primal bound is better than the evaluated objective value of the optimal solution, that is, the primal bound cuts off the best solution.
an unexpected error occurs during the solving process, which is also considered a failure.
All these checks are provided by MIP-DD out of the box and can be used for any other MIP solver as well.
@@
Unlike the debug solution assertions, the validity of these checks is not affected by strong dual reductions or symmetry handling.
@@ 2.3 Workflow
Aggregating modifications into batches can sometimes result in valid minor modifications being overlooked, especially when only a small subset of the elements within a batch is responsible for the solver’s incorrect behavior. Consequently, multiple calls to the same modifier may be required to uncover these hidden valid modifications. This is particularly reasonable as the batch size typically decreases along with the problem size. To address this, MIP-DD uses a stage-round mechanism to call the modifiers. This approach is similar to the workflow of presolving, see, e.g., [19, 20]. In the -th stage only the modifiers with priority at most are called in sequence. This is repeated over several rounds until no more changes are applied. Then, the stage number is incremented or the algorithm terminates if there are no more modifiers to add to the stage cycle. Hence, if in one round the problem remains the same, the next stage is entered to include an additional modifier. Otherwise, the next round is entered to repeat the process with the same set of modifiers.
@@ 2.4 General Recommendations for Applying MIP-DD
Determine batch size.
@@
which is maximum number of solves per modifier invocation.
This way, every successful solve invocation will simplify a certain chunk of the problem at once.
As a rule of thumb, in cases when reproducing a bug takes longer, parameter nbatches
should be set to a lower value.
This can help to reduce the total runtime of MIP-DD.
The final instance with a possibly reduced solving time can then be processed faster with a larger number of batches in order to find further reductions.
Define limits for solves.
@@
Suppress known issues.
By default, all types of fails are treated as bugs.
However, it can happen that, for example, a dual fail turns into a primal fail during the reduction process.
Although discovering further issues is a beneficial side effect, it is sometimes desired to avoid particular types of fails in order to suppress unresolved known issues or prioritize certain bug fixes.
@@
Negative values are reserved for solver-internal errors and positive values encode issues detected by MIP-DD.
Exact solution.
For numerically sensitive issues it is especially important to ensure that the reference solution has feasibility violations as small as possible. To illustrate potential confusion due to an inaccurate reference solution, we consider the simple example
@@ 3 Case Study
The primary objective of Delta debugging is to enable rapid bug tracking and save time during the debugging process. While these time savings are difficult to quantify directly, we use five case studies in this section to demonstrate that MIP-DD often produces significant reductions leading to an effective acceleration and motivation in the debugging process. During the development of the SCIP versions 8.0.1 to 8.1.1, preliminary versions of MIP-DD were applied to significantly reduce the reproducing instances of 24 among all the 51 MIP-related and finally fixed bugs listed in the release reports. Further statistics on the bugs can be found in the repository [7]. Moreover, reduced instances generated by these bugs have been added to the set of instances that are tested regularly for sustainability during SCIP development. They can be found in the SCIP repository222https://github.com/scipopt/scip/tree/master/check/instances/Issue.
Before presenting specific examples, we need to briefly discuss the fact that MIP solvers typically employ floating-point arithmetic. In floating-point arithmetic, a limited amount of memory is allocated to represent a number. Hence, not all numbers can be represented accurately. If more digits are needed than allocated the last digits are omitted leading to a loss in precision. Among further problems, calculation in floating-point arithmetic is neither associative nor distributive meaning that is not necessarily equal to and may yield a different result than . For further information about floating-point arithmetic, we refer to [24].
@@ 3.1 Modifying the Settings to Detect Wrong Calculation of Activities in DualInfer
We first want to highlight the ability of MIP-DD to identify faulty components by manipulating the settings.
For this, we take a look at an instance reported by a SCIP user that has been in the backlog for over three years (SCIP git hash d4c0ad7644
Using the so-called aggressive presolve settings leads to an incorrect solution while default settings result in the correct solution.
@@
In fact, the reason for the incorrect solution is that Dual-Infer incorrectly calculated the min-max-residuum for the subset where represents the index set of all variables and . If applied correctly, for a constraint with the maximal activity is calculated by
@@
the maximum residuum for the first constraint is .
@@
The reason is an index shift where instead of and , and were always applied.
@@ 3.2 Fixing Invalid Separation of Interior Solutions
In this section, we want to analyze a bug that appeared on the MIPLIB [25] instance rococoC10-001000
@@
For this problem with a known true optimal solution value of 11460, a solution with value 23706 was declared optimal after 5776 LP iterations. Applying delta debugging with disabled separators, heuristics, and presolvers as target settings, as well as suppressed primal fails, and 1,000 batches, led to the small integer program
in less than an hour. The small program has the same optimal value as the original problem yet SCIP returns an “optimal” solution of value 12087 after 4 LP iterations. The only remaining non-disabling settings are presolving/maxrounds = -1, constraints/linear/maxprerounds = -1, separating/closecuts/freq = 0, and separating/intobj/freq = 0. This means that apart from some presolving only separators closecuts and intobj are required to reproduce this issue. As closecuts is known as a meta separator, these settings indicated that there might be an issue in separator intobj when called by separator closecuts [26, 27, 28]. Separator closecuts determined an interior point of the LP relaxation which is handed to intobj to be separated. Separator intobj relies on the integrality of the objective function, which is exploited by cutting off a solution with a fractional objective value. This is done by rounding this value up and using it as left-hand side of the objective row. The approach implicitly assumes that there is no feasible solution with a smaller objective value than the infeasible solution to be separated since all undiscovered better solutions would otherwise be cut off. However, for the arbitrary LP-interior points computed by separator closecuts, this does not hold in general. It turned out that the value of the separated solution is approximately 11646.03 indeed exceeding the optimal value 11460. Restricting intobj to optimal solutions of LP relaxations finally resolved this issue.
During preliminary experiments, we observed on this issue that using default settings as target settings, results in a larger final instance with 1528 variables and 301 constraints. The runtime was even comparable to the previously presented setup although disabling plugins usually leads to larger solving times of the single solves. An explanation for this behaviour might be given by the complexity of the solver under default settings which under certain circumstances requires a more complex external problem to reproduce the failing internal state. Nevertheless, targeting default settings is desirable for the generation of lightweight but flexible regular tests. And for other issues, comparably small reproducing instances could be found.